OpenStudy (anonymous):

You have 10 fewer quarters than dimes and 5 fewer nickels than quarters. The total value of coins is 4.75. How many of each denomination do you have?

OpenStudy (anonymous):

d-10=q 10 fewer quarters than dimes q-5=n 5 fewer nickels than quarters n+d+q=4.75 The total value of coins is 4.75 (q-5)+(q+10)+q=4.75 Solve for q, then get dimes and nickles.